OpenStudy (anonymous):

period of sin theta+sin 2 theta/cos theta+cos 2 theta

OpenStudy (anonymous):

sin theta=90

OpenStudy (anonymous):

sin 2 theta=45

OpenStudy (anonymous):

90+45/360+180

OpenStudy (anonymous):

\[\sin \theta+\cos 2 \theta/(\cos \theta+\cos 2\theta)\]

OpenStudy (anonymous):

sorry in numerator sin 2theta

OpenStudy (anonymous):

i thin sin x=2pi

OpenStudy (crashonce):

the period of sinx is 2pi

OpenStudy (anonymous):

so sin 2x=180

OpenStudy (anonymous):

180+360/360+180=1

OpenStudy (crashonce):

@dg2 just saying sin2x = 2sinxcosx

OpenStudy (anonymous):

2*360*360

ganeshie8 (ganeshie8):

haha wish we could do all those crimes :D

OpenStudy (crashonce):

^

OpenStudy (anonymous):

sin 2x=180

OpenStudy (anonymous):

\[\sin \theta +\sin 2 \theta /(\cos \theta+\cos 2\theta)\]

OpenStudy (anonymous):

is it correct?

OpenStudy (anonymous):

i got 1 as answer

Parth (parthkohli):

\[\sin\theta + \sin2\theta = 2\sin(3\theta/2)\cos(\theta/2)\]\[\cos\theta + \cos2\theta = 2\cos(3\theta/2)\cos(\theta/2)\]Divide equations one and two,\[\cdots = \tan(3\theta/2)\]

Parth (parthkohli):

The period of \(\tan(3\theta/2)\) can be easily determined.

OpenStudy (anonymous):

sin 2x=2sinxcosx

OpenStudy (anonymous):

2sinxcosx+sinx

OpenStudy (anonymous):

how 3theta/2 in sin and cos @ParthKohli

Parth (parthkohli):

I used these identities:\[\sin C + \sin D = 2\sin\left(\dfrac{C+D}{2}\right)\cos\left(\dfrac{C-D}{2}\right)\]\[\cos C + \cos D = 2\cos\left(\dfrac{C + D}{2}\right)\cos\left(\dfrac{C - D}{2}\right)\]

OpenStudy (anonymous):

cos2x=2cos^2x-1=1-2sin^2x=cos^2x-sin^2x

OpenStudy (anonymous):

ohh ok how to determine tan 3teta/2

Parth (parthkohli):

OK, so you understood how I got \(\tan(3\theta/2)\) right? I simplified your function to get \(\tan(3\theta/2)\). Do you know what the period of this is?

OpenStudy (anonymous):

yes.tan 3a/2 formula?

Parth (parthkohli):

Do you know what the period of \(\tan\) function is?

OpenStudy (anonymous):

180

OpenStudy (anonymous):

30

OpenStudy (anonymous):

3*180/2

OpenStudy (anonymous):

90/3

Parth (parthkohli):

That's right, 180. So \(\tan(3\theta/2)\) has period = 180 when \(3\theta/2 = 180\implies \theta = 2/3 \times 180 \).

OpenStudy (anonymous):

120

OpenStudy (anonymous):

120 or 30